# Adding Multi-modal Controls to Whole-body Human Motion Generation

### v1

@@ Conclusion
In this paper, we propose ControlMM, a unified framework for whole-body human motion generation with plug-and-play multi-modal control that generalizes across different generative distributions and efficiently handles control signals of varying granularity. Our core design is ControlMM-Attn, which effectively learns and transfers motion knowledge across different distributions by parallel modeling of static and dynamic human topology graphs. ControlMM employs a coarse-to-fine training strategy that achieves fine-grained, plug-and-play control for control signals of varying low-level granularity without the optimization burden of mixed training. For a comprehensive evaluation, we introduce ControlMM-Bench, the first publicly available multi-modal whole-body human motion generation benchmark based on the unified whole-body SMPL-X representation. Extensive experiments show that ControlMM achieves competitive performance across various motion generation tasks compared to state-of-the-art baselines.

### v2

@@ Conclusion
In this paper, we propose ControlMM, a unified framework for whole-body human motion generation with plug-and-play multimodal control that generalizes across different generative distributions and efficiently handles control signals of varying granularity. Our core design is ControlMM-Attn, which effectively learns and transfers motion knowledge across different distributions by parallel modeling of static and dynamic human topology graphs. ControlMM employs a coarse-to-fine training strategy that achieves fine-grained, plug-and-play control for control signals of varying low-level granularity without the optimization burden of mixed training. For a comprehensive evaluation, we introduce ControlMM-Bench, the first publicly available multimodal whole-body human motion generation benchmark based on the unified whole-body SMPL-X representation. Extensive experiments show that ControlMM achieves competitive performance across various motion generation tasks compared to state-of-the-art baselines.